In 2011, a man was 40 and his daughter was 12. In what year was the man just twice his daughter's age?
Answer by ankor@dixie-net.com(22740)

n 2011, a man was 40 and his daughter was 12.
In what year was the man just twice his daughter's age?
:
t = no. of years for this to be true
:
40 + t = 2(12+t)
40 + t = 24 + 2t
40 - 24 = 2t - t
16 = t
:
2011 + 16 = 2027, this will be true
:
:
See if this is true
40+16 = 2(12+16)
56 = 2(28)